OpenStudy (anonymous):

find the equation of the line containing the given pair ofpoints. (2,5) and (6,8)

OpenStudy (anonymous):

m=(8-5)/(6-2)=3/4 y-5=(3/4)(x-2) y-5=(3/4)x-(3/2) y=(3/4)x+(7/2)